# Babysit A Language Model From Scratch: Interactive Language Learning by Trials and Demonstrations

### v1

@@ 4.4 Practice makes perfect in trials
Finally, we conduct experiments to underscore the significance of the student’s active trials in the process of word learning. A student model can learn a word from 3 sources: its own trials, teacher demonstrations, and exposure to the corpus. To determine which source contributes most significantly to learning, we begin by plotting the per-word learning curves alongside the cumulative frequency of word encounters in trials, demonstrations, and the corpus. We observe, interestingly, that the learning curves for certain words exhibit a pronounced correlation with the frequency of these words in trials, as exampled in Figure 9 and 11.
We speculate that this pattern may be associated with the part of speech (POS) of the word. Following Portelance et al. [45], we delve into this phenomenon by focusing on specific subsets of words, including nouns, predicates, and functional words, to explore the relationship further. To evaluate the impact of each source of word acquisition, we consider the cumulative word frequency as a predictor of the word surprisal, and carry out linear regressions complemented by likelihood ratio tests to determine the beta weights for each predictor. Upon analysis, we identify significant collinearity between word frequency in teacher demonstrations and the corpus, as indicated by a high variance inflation factor (VIF), while the word frequency in student trials is less intertwined, exhibiting a moderate VIF below 10. We thus calculate the beta weights separately for pairs (trial, demo) and (trial, corpus), then compute the average beta weights for trials. Together with the Pearson correlation, we summarize the results in Table 3. Negative beta weights signify a negative correlation, with a larger absolute value denoting a stronger association and contribution. Our analysis reveals that the cumulative frequency of words encountered in trials plays a significant role in the acquisition of functional words and predicates. However, this significant contribution does not extend to nouns, indicating a potential impact of active trials on different POS within the learning process. This finding is linguistically intuitive, as function words and predicates are words that require other dependent words to fully express their meaning [23] and thus require more practice. We posit that grounding language in the non-linguistic world is essential for acquiring the meanings of words, particularly for concrete noun [32].
@@ 5.1 Interaction in neural language learning
Researchers have emphasized the role of interaction in computational models of language [6, 66]. Preliminary efforts have been conducted under specific constraints, such as in domain-specific scenarios [47, 73, 4, 61, 34] or considering particular types of dialogue acts [77, 75]. More recently, a series of studies have approached language acquisition through the lens of multimodal referential games [29, 79, 31], emphasizing the importance of pragmatic inference and communicative feedback in speaker-listener interactions. Early work investigated scenarios where the teacher models actively select training data to optimally assist a passive student learner [62]. Nikolaus and Fourtassi [37] adopted setups where student models learn by producing language and receive feedback of communicative success using the BLEU score [42]. However, these models do not receive teachers’ demonstrations or corrections. Our work diverges from these studies as we focus on symmetric teacher-student interactions that require language production from the student and explicit corrective feedback from the teacher, which goes beyond the simple speaker-listener roles. We also adopt massive corpora to study modern transformer-based generative language models, a general approach without domain restrictions. Recent advancements in reinforcement learning from human feedback (RLHF) [40] mark a breakthrough in interactive language learning. This work leverages the implementation infrastructure of RLHF, but diverges in significant ways. Whereas RLHF aims to align a pre-trained language model with human preferences, our objective is to “babysit” a language model from scratch. We refer to existing surveys [27, 78] and discuss RLHF further in Appendix A.1.
Figure 10: A natural example of corrective feedback in child-parent conversation, selected from the CHILDES corpus [33].
@@ 5.2 Interaction in human language learning
To unveil the role of social interactions in language acquisition, a prominent body of research has focused on pragmatic inference – that children exhibit the ability to refine their linguistic knowledge by inferring the communicative intents of others [56, 76, 9]. Besides children’s pragmatic inference about the speakers, many researchers have argued for the importance of caregivers’ feedback to the language produced by children [70], in the form of descriptions, explanations, corrections, etc, to human language development. For example, extensive efforts have been made to examine the effects of communicative feedback on language acquisition, both in developmental psychology [3, 60, 38] and computational modeling [37, 31]. This type of feedback emphasizes the explicit negotiation of mutual understanding with the conversational partner to achieve and maintain common ground [17]. While communicative feedback emphasizes the success and failure of communication, the feedback we study in this study is more akin to corrective feedback. As is shown in Figure 10, this type of feedback involves responses from caregivers, which offer corrections to possible errors in children’s speech, including variants such as negative evidence, reformulations, or recasts [21, 51, 15, 52, 26]. Although the role of corrective feedback has been very well confirmed in second language acquisition [18, 19, 7], researchers largely dispute its availability and effectiveness in human first language acquisition [11, 35]. While our findings suggest that corrective feedback through demonstrations can enhance efficient neural language learning, these results should not be generalized to human language learning, where demonstrations are much less frequent in the noisy feedback that children typically receive.
5.3 Psychologically motivated analysis of language models
Although language models and human language learners differ in their inductive biases and data sources [2, 71], several works have looked into the learnability, proficiency, and efficiency of neural language learning, for example, the relationships between word surprisal in language models to various psycholinguistic variables [45, 32]. Recent efforts have shifted towards exploring the developmental trajectories of language models, rather than their end performance [55, 8, 5, 74], sparkling further investigations into the developmental aspects of psycholinguistics using computational approaches [13, 14, 20]. The scientific rationale behind this is that these models can serve as hypothesis generators or proofs of concept, verifying mechanisms that are practically effective for machines and potentially feasible for human learners [43, 44]. We echo that the benchmark outcomes of language models themselves are insufficient [2, 43], and researchers need to control the factors that may have contributed to the model’s learning process, particularly through conducting ablation studies [71]. In this study, we follow this spirit and investigate the role of trials and demonstrations through systematic computational experiments, assessing its role in neural word acquisition.
@@ 6 Conclusion
Our study is limited by the use of a single reward model focused on corrective feedback. More realistic scenarios should also encompass communicative feedback, with the success of communication serving as a reward. Additionally, Thorndike [63] proposed the idea that a child might instinctively feel satisfaction from producing a sound that echoes a meaningful memory. The design of such an intrinsic reward model is cognitively intriguing, and could aid in scaling student models without an external reward model, offering potential benefits for engineering applications.
@@
One limitation of our approach is the reliance on the Byte Pair Encoding (BPE) tokenizer [57] inherited from GPT-2. Ideally, the method should be tokenizer-free to facilitate the learning of early language elements such as sound effects and animal sounds, e.g., “baa-baa” in CDI [22], which can be crucial for a more natural and foundational language acquisition process.
@@
Acknowledgments and Disclosure of Funding
This work was supported in part by NSF IIS-1949634 and NSF SES-2128623. The authors extend their special appreciation to Susan Gelman and Freda Shi for their valuable feedback (both communicatively and correctively). The authors would like to thank Run Peng, Yichi Zhang, Jacob Sansom, Zheyuan Zhang, Xuejun Zhang for proofreading and their helpful input during discussions.
@@ Appendix A Trials-and-Demonstrations (TnD) Details and Discussions
A.1 Relationship with RLHF
Reinforcement learning enables language systems to learn from feedback in the form of rewards from games [36, 25] or heuristic scores [49, 37]. Recent advancements in reinforcement learning from human feedback (RLHF) have generated considerable excitement, especially in its application to large language models such as ChatGPT [39]. Reinforcement learning is employed to align the model’s policy with human preferences, utilizing human-annotated preference data [81, 40] or AI models acting as proxies for human judgment [1, 30]. We refer to Kaufmann et al. [27] and Zheng et al. [78] for more details. Our work leverages the implementation infrastructure of RLHF but diverges in significant ways: whereas RLHF aims to align an existing language model with human preferences, our objective is to “babysit” a language model from scratch using reinforcement learning, specifically to model the process of receiving and integrating corrective feedback.
A.2 Preliminaries
Inspired by recent work in RLHF, we use Proximal Policy Optimization (PPO) [54] to train the student model. Consider a model $\pi_{\theta}$ whose the current state $s_{i}$ is a sequence of tokens $s_{i}={t_{1},\cdots,t_{i}}$, and receives a reward $r_{i}$. We outline the key components and refer to Zheng et al. [78] for more details.
@@ A.2 Preliminaries
The advantage function $A_{i}$ at step $i$ is estimated by the Generalized Advantage Estimation (GAE) algorithm [53] for a balanced bias-variance trade-off:
@@ B.1 Reproducibility
We include the list of words and the evaluation datasets within the code. [^2] Our code and data are publicly available at https://github.com/sled-group/TnD.
@@
We randomly initialize the GPT-2 student models. For each combination of corpus and baseline, the training process is conducted 5 times for 10k steps, each with a different random seed. We utilize the top-$k$ decoding strategy for language generation, setting $k=20$. The learning rate for reinforcement learning is set to $2e^{-5}$. We follow the default setting for other PPO hyper-parameters, such as a clip range ($\epsilon=0.2$) from the TRL library. All other hyper-parameters for causal language modeling, including the learning rate set at $1e^{-4}$ and a batch size of 128, remain consistent with the training setup of GPT-2 by Chang and Bergen [13].
@@
In TnD training, each experiment to train a student model using TnD is conducted on 2 A40 GPUs for 36 hours. In teacher model pre-training, we distribute the computation over 4 A40 GPUs with batch size 32 per device for 20 hours. To fine-tune the neural age predictor, we use a LLaMA-2-7B model [65] with regression head on our developmental trajectory (text and step pairs) dataset. To ensure the prediction quality and save computation resources, we fine-tune our model using mixed precision on all parameters. The training is distributed over 8 A40 GPUs with batch size 8 per device with fully-shared data parallel. The learning rate is $5e^{-5}$.
B.2 Additional Results
nAoA distributions.
Additional per-word learning curves.
@@ B.2 Additional Results
B.3 Downstream evaluation on NLU tasks
We evaluate the final model on downstream natural language understanding (NLU) tasks. Specifically, we fine-tune the final CLM and TnD models (from the BabyLM corpus) on the BabyLM round 1 NLU evaluation set, which is based on (Super)GLUE [68, 69]. Table 4 shows that the TnD model performs on par with the CLM model, slightly better overall. We observe that the TnD model did significantly better on the Recognizing Textual Entailment (RTE, which requires determining inferential relationships between hypothesis and premise) task, but underperforms on the Question-Answering NLI (QNLI, which requires comprehending longer paragraphs) task.
@@ B.3 Downstream evaluation on NLU tasks
B.4 Discussion and experiment on other possible baselines

### v2

@@ 4.2 Corrective feedback helps knowledge distillation for smaller student models
The current age-based reward design assumes that the student and teacher models are the same size. This section investigates whether such a reward function could distill linguistic knowledge to smaller student models. The original student GPT-2 model has a dimension of $d=768$ (12 attention heads each with a dimension of 64). We now keep all experimental setups untouched but smaller student models with dimensions of 588 ($12\times 49$), 360 ($10\times 36$), and 250 ($10\times 25$) respectively. Figure 6 shows that such efficient language learning can still be observed, even when the setting is translated to smaller models. Each TnD model outperforms the CLM baseline of the same size and even surpasses CLM baselines of large capacity in early steps.
@@ 4.3 Teacher’s word preferences in demonstrations affect student
To explore how the teacher model’s word selection impacts students’ language development, we repeat the experiments where a chosen set of 40 words for each test vocabulary is excluded from teacher demonstrations. To ensure fluent generation, we maintain the presence of essential functional words so these words don’t appear in the 40 chosen words. During the language generation process by the teacher model, if a word from this set is to be decoded, we select the next best alternative, ensuring these words were never presented in teacher demonstrations. We show the learning curves for these excluded words in Figure 7 and present the nAoA in Table 2 (in Appendix). Our findings indicate that the teacher model’s word choices significantly influence the efficiency of word acquisition by the student model. The absence of words from teacher demonstrations leads to slower learning speed for student models, as evidenced by a higher nAoA, although the student models are ultimately able to learn these words from the corpus and their trials.
(a) The word “have.”
(b) The word “now.”
@@ 4.4 Practice makes perfect in trials
Finally, we conduct experiments to underscore the significance of the student’s active trials in the process of word learning. A student model can learn a word from 3 sources: its own trials, teacher demonstrations, and exposure to the corpus. To determine which source contributes most significantly to learning, we begin by plotting the per-word learning curves alongside the cumulative frequency of word encounters in trials, demonstrations, and the corpus. We observe, interestingly, that the learning curves for certain words exhibit a pronounced correlation with the frequency of these words in trials, as exampled in Figure 8 and 9.
We speculate that this pattern may be associated with the part of speech (POS) of the word. Following Portelance et al. (2020), we delve into this phenomenon by focusing on specific subsets of words, including nouns, predicates, and functional words, to explore the relationship further. To evaluate the impact of each source of word acquisition, we consider the cumulative word frequency as a predictor of the word surprisal, and carry out linear regressions complemented by likelihood ratio tests to determine the beta weights for each predictor. Upon analysis, we identify significant collinearity between word frequency in teacher demonstrations and the corpus, as indicated by a high variance inflation factor (VIF), while the word frequency in student trials is less intertwined, exhibiting a moderate VIF below 10. We thus calculate the beta weights separately for pairs (trial, demo) and (trial, corpus), then compute the average beta weights for trials. Together with the Pearson correlation, we summarize the results in Table 3. Negative beta weights signify a negative correlation, with a larger absolute value denoting a stronger association and contribution. Our analysis reveals that the cumulative frequency of words encountered in trials plays a significant role in the acquisition of functional words and predicates. However, this significant contribution does not extend to nouns, indicating a potential impact of active trials on different POS within the learning process. This finding is linguistically intuitive, as function words and predicates are words that require other dependent words to fully express their meaning (Gleitman, 1990) and thus require more practice. We posit that grounding language in the non-linguistic world is essential for acquiring the meanings of words, particularly for concrete noun (Ma et al., 2023).
@@ 5.1 Interaction in neural language learning
Researchers have emphasized the role of interaction in computational models of language (Bisk et al., 2020; Tsuji et al., 2021). Preliminary efforts have been conducted under specific constraints, such as in domain-specific scenarios (Qu and Chai, 2010; Weston, 2016; Bianchi et al., 2021; Stein et al., 2021; Madotto et al., 2021) or considering particular types of dialogue acts (Zhang et al., 2018; Yuan et al., 2019). More recently, a series of studies have approached language acquisition through the lens of multimodal referential games (Lazaridou et al., 2016; Zhu et al., 2022; Liu et al., 2022), emphasizing the importance of pragmatic inference and communicative feedback in speaker-listener interactions. Early work investigated scenarios where the teacher models actively select training data to optimally assist a passive student learner (ter Hoeve et al., 2022). Nikolaus and Fourtassi (2021) adopted setups where student models learn by producing language and receive feedback of communicative success using the BLEU score (Papineni et al., 2002). However, these models do not receive teachers’ demonstrations or corrections. Our work diverges from these studies as we focus on symmetric teacher-student interactions that require language production from the student and explicit corrective feedback from the teacher, which goes beyond the simple speaker-listener roles. We also adopt massive corpora to study modern transformer-based generative language models, a general approach without domain restrictions. Recent advancements in reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) mark a breakthrough in interactive language learning. This work leverages the implementation infrastructure of RLHF, but diverges in significant ways. Whereas RLHF aims to align a pre-trained language model with human preferences, our objective is to “babysit” a language model from scratch. We refer to existing surveys (Kaufmann et al., 2023; Zheng et al., 2023) and discuss RLHF further in Appendix A.2.
5.2 Psychologically motivated analysis of language models
Although language models and human language learners differ in their inductive biases and data sources (Baroni, 2022; Warstadt and Bowman, 2022), several works have looked into the learnability, proficiency, and efficiency of neural language learning, for example, the relationships between word surprisal in language models to various psycholinguistic variables (Portelance et al., 2020; Ma et al., 2023). Recent efforts have shifted towards exploring the developmental trajectories of language models, rather than their end performance (Sellam et al., 2021; Blevins et al., 2022; Biderman et al., 2023; Xia et al., 2023), sparkling further investigations into the developmental aspects of psycholinguistics using computational approaches (Chang and Bergen, 2022; Chang et al., 2023; Evanson et al., 2023). The scientific rationale behind this is that these models can serve as hypothesis generators or proofs of concept, verifying mechanisms that are practically effective for machines and potentially feasible for human learners (Portelance, 2022; Portelance and Jasbi, 2023). We echo that the benchmark outcomes of language models themselves are insufficient Baroni (2022); Portelance (2022), and researchers need to control the factors that may have contributed to the model’s learning process, particularly through conducting ablation studies Warstadt and Bowman (2022). In this study, we follow this spirit and investigate the role of trials and demonstrations through systematic computational experiments, assessing its role in neural word acquisition.
@@ 5.3 Interaction in human language learning
Social interactions are crucial in language acquisition, and the role of caregivers’ feedback has been extensively explored in the field of developmental psychology (Bates et al., 1975; Saxton et al., 2005; Warlaumont et al., 2014). While our findings suggest that corrective feedback could enhance efficient neural word learning, these results should not be generalized to human language learning. We discuss more about the role of interaction in human language learning in Appendix A.1.
@@ 6 Conclusion
Additional experiments.
Due to the limited space and primary research scope of this work, we present experiments about other possible baselines besides controllable ablation, downstream natural language understanding task performances, and the robustness of the findings under different hyper-parameters in Appendix B.
@@
Our study is limited by the use of a single reward model focused on corrective feedback. More realistic scenarios should also encompass communicative feedback, with the success of communication serving as a reward. Additionally, Thorndike (1911) proposed the idea that a child might instinctively feel satisfaction from producing a sound that echoes a meaningful memory. The design of such an intrinsic reward model is cognitively intriguing, and could aid in scaling student models without an external reward model, offering potential benefits for engineering applications.
@@
One limitation of our approach is the reliance on the Byte Pair Encoding (BPE) tokenizer (Sennrich et al., 2016) inherited from GPT-2. Ideally, the method should be tokenizer-free to facilitate the learning of early language elements such as sound effects and animal sounds, e.g., “baa-baa” in CDI (Fenson et al., 2006), which can be crucial for a more natural and foundational language acquisition process.
@@
Acknowledgement
This work was supported in part by NSF IIS-1949634 and NSF SES-2128623. Ziqiao Ma is supported in part by the Weinberg Cognitive Science Fellowship. The authors extend their special appreciation to Susan Gelman and Freda Shi for their valuable feedback. The authors would like to thank Run Peng, Yichi Zhang, Jacob Sansom, Zheyuan Zhang, Xuejun Zhang for proofreading and their helpful input during discussions. We thank all anonymous reviewers for their feedback.
@@ Appendix A Trials-and-Demonstrations (TnD) Details and Discussions
A.1 Interaction in human language learning
To unveil the role of social interactions in language acquisition, a prominent body of research has focused on pragmatic inference – that children exhibit the ability to refine their linguistic knowledge by inferring the communicative intents of others (Senju and Csibra, 2008; Yurovsky and Frank, 2017; Bohn and Frank, 2019). In addition, other researchers have argued for the importance of caregivers’ feedback to the language produced by children (Warlaumont et al., 2014), in the form of descriptions, explanations, corrections, etc, to human language development. For example, extensive efforts have been made to examine the effects of communicative feedback on language acquisition, both in developmental psychology (Bates et al., 1975; Snow et al., 1996; Nikolaus and Fourtassi, 2023) and computational modeling (Nikolaus and Fourtassi, 2021; Liu et al., 2022). This type of feedback emphasizes the explicit negotiation of mutual understanding with the conversational partner to achieve and maintain common ground Clark (1996). While communicative feedback emphasizes the success and failure of communication, the feedback we study in this study is more akin to corrective feedback. This type of feedback involves responses from caregivers, which offer corrections to possible errors in children’s speech, including variants such as negative evidence, reformulations, or recasts (Farrar, 1992; Saxton, 2000; Chouinard and Clark, 2003; Saxton et al., 2005; Hiller, 2016). Although corrective feedback is shown to be helpful in second language acquisition (El Tatawy, 2002; Ellis et al., 2006; Bitchener and Ferris, 2011), researchers largely dispute its availability and effectiveness in human first language acquisition (Brown, 1970; Marcus, 1993). While our findings suggest that corrective feedback through demonstrations can enhance efficient neural language learning, these results should not be generalized to human language learning, where demonstrations are much less frequent in the noisy feedback that children typically receive.
A.2 Relationship with RLHF
Reinforcement learning enables language systems to learn from feedback in the form of rewards from games (Narasimhan et al., 2015; He et al., 2016) or heuristic scores (Ranzato et al., 2016; Nikolaus and Fourtassi, 2021). Recent advancements in reinforcement learning from human feedback (RLHF) have generated considerable excitement, especially in its application to large language models such as ChatGPT (OpenAI, 2022). Reinforcement learning is employed to align the model’s policy with human preferences, utilizing human-annotated preference data (Ziegler et al., 2019; Ouyang et al., 2022) or AI models acting as proxies for human judgment (Bai et al., 2022; Lee et al., 2023). We refer to Kaufmann et al. (2023) and Zheng et al. (2023) for more details. Our work leverages the implementation infrastructure of RLHF but diverges in significant ways: whereas RLHF aims to align an existing language model with human preferences, our objective is to “babysit” a language model from scratch using reinforcement learning, specifically to model the process of receiving and integrating corrective feedback.
A.3 Preliminaries
Inspired by recent work in RLHF, we use Proximal Policy Optimization (PPO) Schulman et al. (2017) to train the student model. Consider a model $\pi_{\theta}$ whose the current state $s_{i}$ is a sequence of tokens $s_{i}={t_{1},\cdots,t_{i}}$, and receives a reward $r_{i}$. We outline the key components and refer to Zheng et al. (2023) for more details.
@@ A.3 Preliminaries
The advantage function $A_{i}$ at step $i$ is estimated by the Generalized Advantage Estimation (GAE) algorithm Schulman et al. (2016) for a balanced bias-variance trade-off:
@@ B.1 Reproducibility
We include the list of words and the evaluation datasets within the code. [^2] Our code and data will be made publicly available upon the acceptance of this work.
@@
We randomly initialize the GPT-2 student models. For each combination of corpus and baseline, the training process is conducted 5 times for 10k steps, each with a different random seed. We utilize the top-$k$ decoding strategy for language generation, setting $k=20$. The learning rate for reinforcement learning is set to $2e^{-5}$. We follow the default setting for other PPO hyper-parameters, such as a clip range ($\epsilon=0.2$) from the TRL library. All other hyper-parameters for causal language modeling, including the learning rate set at $1e^{-4}$ and a batch size of 128, remain consistent with the training setup of GPT-2 by Chang and Bergen (2022).
@@
In TnD training, each experiment to train a student model using TnD is conducted on 2 A40 GPUs for 36 hours. In teacher model pre-training, we distribute the computation over 4 A40 GPUs with batch size 32 per device for 20 hours. To fine-tune the neural age predictor, we use a LLaMA-2-7B model (Touvron et al., 2023) with regression head on our developmental trajectory (text and step pairs) dataset. To ensure the prediction quality and save computation resources, we fine-tune our model using mixed precision on all parameters. The training is distributed over 8 A40 GPUs with batch size 8 per device with fully-shared data parallel. The learning rate is $5e^{-5}$.
@@
B.2 Additional Results
nAoA distributions.
Surprisal and learning curves.
(a) Plateau v.s. unigram surprisal in BookCorpus.
(b) Plateau v.s. unigram surprisal in BabyLM.
Figure 10: Sigmoid plateau v.s. unigram surprisal.
We observe a similar pattern reported by Chang and Bergen (2022) that learning curves tend to level off at a local plateau, which aligns with the unigram surprisal. This phenomenon renders the single-sigmoid model unreliable for capturing the complexity, and we adopt a double-sigmoid function to fit the learning curve. We run regression between the plateau of double-sigmoid curves and the unigram surprisals calculated from all sources of word occurrences (Figure 10). We find a strong correspondence between the plateau and unigram frequency, suggesting that the double-sigmoid function is a better option than the single-sigmoid function to fit learning curves. To analyze longer learning curves of up to 1M steps, more complex functions such as linear GAMs have been adopted (Chang et al., 2023). For our purposes, a double-sigmoid function suffices.
B.3 Teacher’s word preferences in demonstrations affect student
@@
Additional per-word learning curves.
B.4 Downstream evaluation on NLU tasks
We evaluate the final model on downstream natural language understanding (NLU) tasks. Specifically, we fine-tune the final CLM and TnD models (from the BabyLM corpus) on the BabyLM round 1 NLU evaluation set, which is based on (Super)GLUE (Wang et al., 2018, 2019). Table 4 shows that the TnD model performs on par with the CLM model, slightly better overall. We observe that the TnD model did significantly better on the Recognizing Textual Entailment (RTE, which requires determining inferential relationships between hypothesis and premise) task, but underperforms on the Question-Answering NLI (QNLI, which requires comprehending longer paragraphs).
B.5 Discussion and experiment on other possible baselines
Causal language modeling using teacher-generated text (naive content distillation).
[…]
Causal language modeling using teacher-generated text (naive content distillation).
Our main difference from the existing knowledge distillation methods for LLMs is that we train a student model from scratch, rather than requiring a pre-trained but weaker student model. For example, Ko et al. (2024) adopt the T5 language model as the student. Closest to our work is Nikolaus and Fourtassi (2021), which also adopts an ablation study setup. We did not compare to their method as a baseline as they noticed that combining production and perception-based language learning does not work from scratch. To the best of our knowledge, our TnD method is the only interactive language learning algorithm that functions well with a student model from scratch.
Using the corpus sentences as demonstrations.
While it is possible to directly use the ground truth sentences from the training corpus as demonstrations (ter Hoeve et al., 2022), it can be very difficult to adapt the teacher model’s behaviors for our controlled studies. As a result, we use a pre-trained language model as the proxy for demonstration generation, rather than using the original text. In our preliminary experiments, using ground truth sentences from the training corpus as demonstrations do not lead to a noticeable difference from model-generated demonstrations.
B.6 The robustness of results and findings over hyper-parameters